OpenStudy (help!!!!):

one muber is 3 times another number less than 12. If the first number is 108, what is the second number?

OpenStudy (raden):

let the first numer is x and the 2nd is y, so x = 3(y-12) now if given x=108, it means 108 = 3(y-12) solve for y

OpenStudy (raden):

*number

OpenStudy (help!!!!):

ohhh give me a sec to understand

OpenStudy (help!!!!):

120 is the second number

OpenStudy (help!!!!):

but how did you know how you put it in a equation?

OpenStudy (raden):

i just transleting mathematic statement to algebra languange :)

OpenStudy (raden):

its principle algebra, u have to know it

OpenStudy (help!!!!):

thanks

OpenStudy (raden):

welcome... btw, what's ur answer did u get

OpenStudy (help!!!!):

120

OpenStudy (raden):

108 = 3(y-12) divided by 3 to both sides, gives 108/3 = y-12 36 = y-12 add by 12 to both sides 36+12=y-12+12 48=y hope it helped